OpenStudy (anonymous):

The length of a rectangle is 1 less than twice its width, and the area of the rectangle is 55ft^2 . Find the dimensions of the rectangle

OpenStudy (amistre64):

Area = lw l=2w-1 area = (2w-1)(w) 55 = 2w^2 -w 2w^2 -w -55 = 0 this is what i got so far

OpenStudy (amistre64):

w =1/4 +- sqrt(1-(4)(2)(-55))/4 just let me get the calculator... do you know the quadratic formula?

OpenStudy (amistre64):

1/4 +- sqrt(1 + 440)/4 1/4 +- sqrt(441)/4 this better work :)

OpenStudy (anonymous):

so is this the final answer i also need to know the length and width

OpenStudy (amistre64):

1/4 +- 21/4...-21/4 makes no sense, so we can toss it out the window. 1/4 + 21/4 = 22/4 = 11/2 the width of this thing is 11/2 l=2(11)/2-1 = 11-1 = 10 length = 10 width = 11/2 does that make sense? and does it fit?

OpenStudy (amistre64):

55 = (10)(11)/(2) 55 = 110/2 55 = 55 it fits, and its works.... yay!!!

OpenStudy (anonymous):

OMG YOUR AMAZING!!! THANK YOU SO MUCH I JUST FINISHED THE CLASS YOUR AMAZING!!!!!!!!!!!!!!!!!!! AHHHHHHHHHHHHHHHHHHHHH IM JUMPING OUT OF MY CHAIR THANK YOU SO MUCH

OpenStudy (amistre64):

watch out for the ceiling fan :)

OpenStudy (anonymous):

Lol thanks, jocks like me aint built to do math lol

OpenStudy (anonymous):

Thank you so much once again have a great night!

OpenStudy (amistre64):

Ciao :)